# Understanding-in-Generation:Reinforcing Generative Capability of Unified Model via Infusing Understanding into Generation

### v1

@@ 1 Introduction
First, for verification-based reasoning (e.g., ImageCoT Zhang et al. (2025a)), as illustrated in Figure 1 (a), this method constructs multiple generative branches through repeated sampling and assesses the generative potential throughout the generation process. Specifically, the state of each branch is evaluated to determine whether it still has the potential to continue generating. If the assessment gives a negative result, the branch is terminated. In verification-based reasoning, the understanding of the unified model is applied exclusively to the verification of intermediate images, selecting the best output from among numerous samples. However, in this framework, the understanding capability is employed merely as a tool for validation and filtering, rather than providing effective guidance during the generation. As a result, the generative ability remains confined to scopes that can be reached through repeated sampling. As shown in Figure 1 (a), given the input prompt, [”There is a cup positioned behind the woman”], all outputs consistently place the cup in front of the woman. Consequently, even the best result obtained through verification-based reasoning fails to satisfy the spatial relationship specified in the input prompt.
@@
Figure 1: The overall of (a) verify-based reasoning, (b) prompt reasoning, and (c) our UiG.
In this paper, we introduce Understanding-in-Generation, an effective text-to-image reasoning framework for unified models that leverages their strong understanding capabilities to guide and enhance image generation. The ‘out-of-the-box’ insight of our proposed UiG is to integrate generative guidance through the strong understanding capabilities during the reasoning process, thereby addressing the limitations of generative abilities. Furthermore, our UiG demonstrates the promising potential of unified models to enhance generation through understanding, powered by using the “Image Editing” bridge to infuse understanding into generation.
@@ 3.2 The Understanding-in-Generation Reasoning
We present Understanding-in-Generation reasoning in Figure 2. UiG enhances the generative capabilities of unified models by infusing their strong understanding abilities into the step-by-step reasoning process. The core insight of UiG is to leverage the powerful understanding capabilities of unified models to provide effective guidance throughout the generative reasoning steps. Specifically, given an input text prompt, we first generate an initial image $Image_{0}$ using the original generative capacity of the unified model.
@@
$Image_{i}=\text{Editing}\left(\text{Combine}\left\{Image_{i-1},\ Prompt^{\text{edit}}_{i}\right\}\right),$ (5)
@@ 6 Conclusion
In this paper, we propose Understanding-in-Generation, an effective reasoning framework for text-to-image generation, which mitigates the limitation of generative capabilities via infusing the understanding guidance. Our UiG demonstrates the promising potential of unified models to reinforce their generation capabilities through their strong understanding capabilities. We evaluate our reasoning method on both the TIIF and WISE benchmarks, and the experiment results show significant performance improvements over the existing text-to-image reasoning methods.
Limitation and Future Work: Our UiG is not available for video generative models. Therefore, the future work will focus on improving the reasoning for video generation.
Reproducibility Statement: To ensure reproducibility, we provide the full project code and the experiment results on TIIF and WISE benchmarks in the Supplementary Materials.
@@ B.2 More Details of Experiment Setup
All reported experiments in this paper were conducted on NVIDIA H100 GPUs. To ensure fair comparisons, we set the random seed to a fixed value of 42. For further code details, please refer to the complete code provided in our Supplementary Materials.
Appendix C Additional Visual Results
@@ C.2 More Visualization of Reasoning Process
We show additional visualization of the UiG reasoning process in Figure 12 13 14 15. The visualization illustrate the full reasoning process of our proposed UiG, which demonstrates the effectiveness of our UiG to infuse understanding capabilities into the generation step by step.
@@
Appendix D User Study

### v2

@@ 1 Introduction
First, for verification-based reasoning (e.g., ImageCoT Zhang et al. (2025a)), as illustrated in Figure 1 (a), this method constructs multiple generative branches through repeated sampling and assesses the generative potential throughout the generation process. Specifically, the state of each branch is evaluated to determine whether it still has the potential to continue generating. If the assessment gives a negative result, the branch is terminated. Here, the understanding of the unified model is applied exclusively to the verification of intermediate images, selecting the best output from among numerous samples. However, in this framework, the understanding capability is employed merely as a tool for validation and filtering, rather than providing effective guidance during the generation. As a result, the generative ability remains confined to scopes that can be reached through repeated sampling. As shown in Figure 1 (a), given the input prompt, [”There is a cup positioned behind the woman”], all outputs consistently place the cup in front of the woman. Consequently, even the best result obtained through verification-based reasoning fails to satisfy the spatial relationship specified in the input prompt.
@@
Figure 1: The overall of (a) Verification-based reasoning, (b) prompt reasoning, and (c) our UiG.
In this paper, we introduce Understanding-in-Generation, an effective text-to-image reasoning framework for unified models that leverages their strong understanding capabilities to guide and enhance image generation. The ‘out-of-the-box’ insight of our proposed UiG is to integrate generative guidance through the strong understanding capabilities during the reasoning process, thereby addressing the limitations of generative abilities. Meanwhile, our UiG demonstrates the promising potential of unified models to enhance generation through understanding, powered by using the “Image Editing” bridge to infuse understanding into generation.
@@ 3.2 The Understanding-in-Generation Reasoning
We present Understanding-in-Generation reasoning in Figure 2. UiG enhances the generative capabilities of unified models by infusing their strong understanding abilities into the step-by-step reasoning process. The core insight of UiG is to leverage the powerful understanding capabilities of unified models to provide effective guidance throughout the generative reasoning steps. Specifically, given an input text prompt $Prompt$, we first generate an initial image $Image_{0}$ using the original text-to-image generative function $\text{Generate}_{t2i}(\cdot)$ of the unified model.
@@
$Image_{i}=\text{Generate}_{editing}\left(\text{Combine}\left\{Image_{i-1},\ Prompt^{\text{edit}}_{i}\right\}\right),$ (5)
@@ 6 Conclusion
In this paper, we propose Understanding-in-Generation, an effective reasoning framework for text-to-image generation, which mitigates the limitation of generative capabilities via infusing the understanding guidance. Our UiG treats “Image Editing” as the bridge to incorporate understanding into generation, effectively steering the reasoning process in the correct direction and significantly enhancing the generative capability for the unified model. Our UiG demonstrates the promising potential of unified models to reinforce their generation capabilities through their strong understanding capabilities. We evaluate our reasoning method on both the TIIF and WISE benchmarks, and the experiment results show significant performance improvements over the baselines.
Limitation and Future Work: Our UiG is not available for video generative models. Therefore, future work will focus on improving the reasoning for video generation.
@@ B.2 More Details of Experiment Setup
All reported experiments in this paper were conducted on NVIDIA A100 GPUs. To ensure fair comparisons, we set the random seed to a fixed value of 42.
Appendix C Additional Visual Result
@@ C.2 More Visualization of Reasoning Process
We show additional visualization of the UiG reasoning process in Figure 11 12 13 14. The visualization illustrate the full reasoning process of our proposed UiG, which demonstrates the effectiveness of our UiG to infuse understanding capabilities into the generation step by step.
@@
Appendix D Additional Discussion
D.1 Failure Case
Figure 15: The failure cases.
Our UiG demonstrates a significant improvement in the generative capabilities of unified models, leveraging their robust understanding capabilities. However, due to the limitations in the spatial understanding capabilities of the unified model, generating images that depict rare spatial relationships between objects can lead to certain “inconsistencies”. As shown in Figure 15 (a), while the UiG successfully generates the correct positional relationship between the giraffe and the airplane (the giraffe is behind the airplane) compared to the base model, the relative size relationship between the giraffe and the airplane appears inconsistent. Similarly, as illustrated in Figure 15 (b), compared to the base model, the UiG still generates the image that aligns with the text prompt. However, the keys behind the little girl remain in a floating” state, defying physical laws. This limitation will be addressed as the spatial understanding capabilities of the base model improve. Future work will focus on enhancing the coordination of complex spatial relationships.
D.2 Trade-off
Table 5: Analysis of the Trade-off. We analyze the trade-off between performance and latency on the long prompt setting of the TIIF benchmark.
Model Performance $\uparrow$ Average Latency (s) $\downarrow$
Verify-based Reasoning 67.18 391.92
Prompt Reasoning 67.90 108.23
Our UiG 71.11 157.61
As shown in Table 5, we analyze the trade-off between performance and latency across Verification-based reasoning, prompt reasoning, and our proposed UiG. To ensure a fair comparison, the analysis is conducted using the same base model (BAGEL Deng et al. (2025)). Our UiG achieves the best performance with lower latency in inference time compared to Verification-based reasoning, although it lags slightly behind prompt reasoning in terms of performance. These results demonstrate the overall advantage of UiG in balancing performance and latency.
Appendix E User Study